OpenStudy (anonymous):

how do i do this problem? x=9-4y x+4y=11 i need step by step help

OpenStudy (campbell_st):

take the 1st equation and substitute it into the 2nd equation (9 - 4y) + 4y = 11 then solve... interesting solution.

OpenStudy (anonymous):

i would get an answer that looks like this x=99-48y which is an empty solution

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

here condition of no solution will apply where a1/a2=b1/b2 is not equal to c1/c2